# Classification of foliations via disintegration maps

### v1

@@ 4. On foliations and disintegration maps
Consider $X=\{(x_{1},x_{2})\in\mathbb{R}^{2}~:~x_{1}^{2}+x_{2}^{2}\leq R^{2}\}$, $R>0$, $\mu\in\mathscr{P}(X)$ with full support and uniformly distributed, and $Y=(0,R]$. Consider a foliation of $X$ into circles, that is, consider a foliation $\mathcal{F}$ of $X$ such that $F_{y}=\{(x_{1},x_{2})\in\mathbb{R}^{2}~:~x_{1}^{2}+x_{2}^{2}=y^{2}\}$, for each $y\in Y$, and $\pi:X\to Y$ such that $\pi((x_{1},x_{2}))=\sqrt{x_{1}^{2}+x_{2}^{2}}=y$. To simplify the analysis, consider this setup in polar coordinates: consider $x_{1}=r\cos(\theta)$, $x_{2}=r\sin(\theta)$. Using Cartesian coordinates, an infinitesimal area element can be calculated as
$dA=dx_{1}~dx_{2}.$
To convert it to polar coordinates, the Jacobian determinant $J$ of the coordinate conversion formula has to be considered. In this case:
$J(r,\theta)=\begin{vmatrix}\frac{\partial x_{1}}{\partial r}&\frac{\partial x_{1}}{\partial\theta}\\ \frac{\partial x_{2}}{\partial r}&\frac{\partial x_{2}}{\partial\theta}\end{vmatrix}=\begin{vmatrix}\cos(\theta)&-r\sin(\theta)\\ \sin(\theta)&r\cos(\theta)\end{vmatrix}=r.$
Then,
$dA=r~dr~d\theta.$
Since $\mu$ is a uniformly distributed probability measure on $X$, we can write:
@@
Define $\nu:=\pi_{*}\mu$. Note that, give a open interval $(a,b)\subset(0,R]$, $\nu((a,b))$ is the area between circles of radii $a$ and $b$, normalized by $A$:
@@
$\int_{Y}\mu_{y}~\nu=\int_{0}^{R}\int_{0}^{2\pi}\Big{(}\frac{1}{2\pi}\delta_{y}d\theta\Big{)}\frac{2r}{R^{2}}~dr=\frac{1}{\pi R^{2}}\iint r~dr~d\theta$
so we recover the measure $\mu$ by integrating the family $\mu_{y}$ with respect to $\nu$. Moreover, $d_{Y}(y,y^{\prime}):=d(\pi^{-1}(y),\pi^{-1}(y^{\prime}))=|y-y^{\prime}|$. For simplicity, consider $W_{1}$. One way to write the 1-Wasserstein distance is
@@
Then, the disintegration map $\tilde{f}$ of $\mu$ with respect to $\nu$ is a isometry, and $\mathcal{E}_{1}(\tilde{f})=1$.
Similarly, consider the case $X=\{(x_{1},x_{2})\in\mathbb{R}^{2}~:~x_{1}^{2}+\lambda^{2}x_{2}^{2}\leq R^{2}\}$, with $R>0$ and $\lambda=1+\varepsilon$, $\varepsilon$ positive and sufficiently small, $\mu\in\mathscr{P}(X)$ with full support and uniformly distributed, $Y=(0,R]$, and the foliation of $X$ is given into ellipses, that is, consider a foliation $\mathcal{F}$ of $X$ such that $F_{y}=\{(x_{1},x_{2})\in\mathbb{R}^{2}~:~x_{1}^{2}+\lambda^{2}x_{2}^{2}=y^{2}\}$, for each $y\in Y$, and $\pi:X\to Y$ such that $\pi((x_{1},x_{2}))=\sqrt{x_{1}^{2}+\lambda^{2}x_{2}^{2}}=y$. Converting into polar coordinates: consider $x_{1}=r\cos(\theta)$, $x_{2}=\frac{r}{\lambda}\sin(\theta)$, and
$J(r,\theta)=\begin{vmatrix}\frac{\partial x_{1}}{\partial r}&\frac{\partial x_{1}}{\partial\theta}\\ \frac{\partial x_{2}}{\partial r}&\frac{\partial x_{2}}{\partial\theta}\end{vmatrix}=\begin{vmatrix}\cos(\theta)&-r\sin(\theta)\\ \frac{1}{\lambda}\sin(\theta)&\frac{r}{\lambda}\cos(\theta)\end{vmatrix}=\frac{r}{\lambda}.$
Then,
$dA=\frac{r}{\lambda}~dr~d\theta.$
Since $\mu$ is a uniformly distributed probability measure on $X$, we can write:
@@
Define $\nu:=\pi_{*}\mu$. Note that, give a open interval $(a,b)\subset(0,R]$, $\nu((a,b))$ is the area between ellipses of major axis $a$ and $b$, normalized by $A$:
@@
Figure 3. Normalized arc length as a function of $\theta$ for $\lambda$ equals to $1,1.1,1.5$ and $2$.
Therefore, the measures $\mu_{y}$ have a different dependence on $\theta$:
$d\mu_{y}(r,\theta)=\frac{\frac{dL_{y}(\theta)}{d\theta}}{\mathcal{L}_{y}}~\delta_{y}~d\theta$
where
$L_{y}(\theta)=\int_{0}^{\theta}y~\sqrt{1-(1-\lambda^{-2})\cos^{2}(\tilde{\theta})}~d\tilde{\theta}$
is the arc length of the ellipse $x_{1}^{2}+\lambda^{2}x_{2}^{2}=y^{2}$ for angle between 0 and $\theta$, and $\mathcal{L}_{y}=L_{y}(2\pi)$. By the Fundamental Theorem of Calculus
@@
In this case, $d_{Y}(y,y^{\prime})=d(\pi^{-1}(y),\pi^{-1}(y^{\prime}))=\frac{1}{\lambda}|y-y^{\prime}|$ and, to calculate the $1$-Wasserstein distance between $\mu_{y}$ and $\mu_{y^{\prime}}$, we still have the radial transport, although $d(x,T(x))$ depends of $\theta$: for $x=y\cos(\theta)$, $d(x,T(x))=R_{y^{\prime}}(\theta)-R_{y}(\theta)$, where
@@
Therefore, considering $\hat{f}_{\lambda}$ the disintegration map of $\mu$ with respect to $\nu$, we have
$\mathcal{E}_{1}(\hat{f}_{\lambda})=\frac{2\pi y}{\mathcal{L}_{y}},$
that is, the energy is the ratio between the length of the circle with radius $y$ and the length of the ellipse with major axis equal to $y$. Table 1 displays the values of $\mathcal{L}_{y}$ for $\lambda=1$ (circle), $\lambda=1,001$, $\lambda=1,01$, $\lambda=1,1$, $\lambda=1,5$ and $\lambda=2$. Table 2 summarizes the computed values of $\mathcal{E}_{1}(\hat{f}_{\lambda})$ corresponding to the cases in Table 1. Observe that $\mathcal{E}_{1}(\hat{f}_{\lambda})$ is independent of $y$.
@@ 11 $1,000499687$ $1,004968906$ $1,047025412$ $1,188089106$ $1,297046785$
Figure 4 depicts the dependence of $\mathcal{E}_{1}(\hat{f}_{\lambda})$ on $\lambda$. As shown, $\mathcal{E}_{1}$ grows smoothly with increasing foliation perturbation.
Figure 4. $\mathcal{E}_{1}(\hat{f}_{\lambda})$ as a function of $\lambda$
This example highlights the versatility of our approach: beyond classifying foliations on $X$ via disintegration maps, it also provides a framework for analyzing structural perturbations in foliations. Notably, the concepts developed here have potential applications across dynamical systems, machine learning, and stochastic analysis.

### v2

@@ 4. On foliations and disintegration maps
We will show that there exists a disintegration $\{\mu_{y}\}$, such that $|\nabla f(y)|_{p}=1$ for $\nu$-almost every $y$, but the metric measure foliation condition does not hold. Consider X=[2,1]X=[-2,1], $Y=[0,1]$ and $\pi:X\to Y$ such that
$\pi(x)=\begin{cases}x&,\text{if}~x\geq 0\\ (\phi+\text{id})^{-1}(-x)&,\text{if}~x<0\end{cases}$
where $\phi$ is the Cantor ternary function. Consider a measure $\mu$ on $X$ such that $\nu=\pi_{*}\mu$ is a uniform measure on $Y$ and there exists a disintegration of $\mu$ with respect to $\nu$ such that the conditional measures are
$\mu_{y}=\frac{1}{2}\delta_{y}+\frac{1}{2}\delta_{\varphi(y)}$
where $\varphi(y)=\phi(y)+y$. Note that, for $y^{\prime},y^{\prime\prime}\in Y$,
$d(\pi^{-1}(y^{\prime}),\pi^{-1}(y^{\prime\prime}))=$ $d(\{y^{\prime},\varphi(y^{\prime})\},\{y^{\prime\prime},\varphi(y^{\prime\prime})\})$
$=$ $\min\left\{|y^{\prime}-y^{\prime\prime}|,|\varphi(y^{\prime}),\varphi(y^{\prime})|\right\}$
$=$ $|y^{\prime}-y^{\prime\prime}|.$
Moreover, let $\mathcal{U}$ be the union of the maximum intervals where the Cantor function is constant, i.e., $\mathcal{U}=(\frac{1}{3},\frac{2}{3})\cup(\frac{1}{9},\frac{2}{9})\cup(\frac{7}{9},\frac{8}{9})\cup\dots$. Note that $\mathcal{U}$ is dense in $[0,1]$, $\nu(\mathcal{U})=1$, and for $y\in\mathcal{U}$, for all $y^{\prime},y^{\prime\prime}\in B_{\varepsilon}(y)$, $\varepsilon<d_{Y}(y,\mathcal{U}^{c})$, we have $W_{p}(\mu_{y^{\prime}},\mu_{y^{\prime\prime}})=|y^{\prime}-y^{\prime\prime}|$. Therefore, $|\nabla f(y)|_{p}=1$ for $\nu$-almost every $y\in Y$.
Example 4.4 justifies the hypothesis of taking the supremum over all $y\in Y$ in Definition 4.2. If we had defined the energy functional using an essential supremum, for example, pathological cases like this would arise, where the other hypotheses of Theorem A are satisfied, but we do not have a $p$-metric measure foliation. Another important remark: when considering a general disintegration $\{\mu_{y}\}$, which is defined $\nu$-almost everywhere, we could define measures at the missing points $y$, but this does not alter the geometric structure on $X$. Moreover, note that by requiring that the conditional measures have full support, the fiber structure becomes detectable by the energy functional. The following example illustrates a case in which Theorem A does not hold if the full support condition is not assumed.
Example 4.5.
Consider $X=[-1,1]\times\mathbb{R}$, $Y=[0,1]$, $\pi:X\to Y$ such that $\pi^{-1}(y)$ is an ellipse given by $x_{1}^{2}+\lambda^{2}x_{2}^{2}=y^{2}$, $\lambda>1$, with major radius $y$. Consider a measure $\mu\in\mathcal{M}_{+}(X)$ such that the disintegration of $\mu$ with respect to $\nu$ is $\mu_{y}=\delta_{(0,\frac{y}{\lambda})}$. In this case,
$d(\pi^{-1}(y),\pi^{-1}(y^{\prime}))=\left|\frac{y}{\lambda}-\frac{y^{\prime}}{\lambda}\right|=W_{p}(\mu_{y},\mu_{y^{\prime}})$
for every $y,y^{\prime}\in Y$, but $\{\pi^{-1}(y)\}$ does not define a metric foliation. A representation of the case $\lambda=1.5.$ is provided in Figure 2.
Figure 2. Illustration for the case in which the conditional measures do not have full support: $d(\pi^{-1}(y),\pi^{-1}(y^{\prime}))=W_{p}(\mu_{y^{\prime}},\mu_{y^{\prime\prime}})$ for every $y^{\prime},y^{\prime\prime}\in Y$, but $\{\pi^{-1}(y)\}$ (the ellipses) are not a metric foliation.
In the next example, we analyze the disintegration of a measure $\mu$ under a flow $\varphi$. For $t=0$, we have the characterization of a metric measure foliation given by Theorem A. We use the energy functional to analyze the change in the foliation associated with the measure $\mu_{t}$, which is the push-forward of $\mu$ by the flow $\varphi$ at time $t$. This example shows that the energy functional exhibits a certain sensitivity to smooth changes in the foliation structure.
Example 4.6.
Consider $X=\{(x_{1},x_{2})\in\mathbb{R}^{2}~:~x_{1}^{2}+x_{2}^{2}\leq R^{2}\}$, $R>0$, $\mu\in\mathscr{P}(X)$ with full support and uniformly distributed on $X$, and $Y=(0,R]$. Consider a foliation of $X$ into circles, that is, consider a foliation $\mathcal{F}$ of $X$ such that $F_{y}=\{(x_{1},x_{2})\in\mathbb{R}^{2}~:~x_{1}^{2}+x_{2}^{2}=y^{2}\}$, for each $y\in Y$, and $\pi:X\to Y$ such that $\pi((x_{1},x_{2}))=\sqrt{x_{1}^{2}+x_{2}^{2}}=y$. Let $\tilde{f}$ be the disintegration map of $\mu$ with respect to $\pi_{*}\mu$. We will show that $\mathcal{E}_{1}(\tilde{f})=1$.
Consider this setup in polar coordinates. Since $\mu$ is a uniformly distributed probability measure on $X$, we can write:
@@
Define $\nu:=\pi_{*}\mu$. Note that, given a open interval $(a,b)\subset(0,R]$, $\nu((a,b))$ is the area between circles of radii $a$ and $b$, normalized by $A$:
@@
$\int_{Y}\mu_{y}~d\nu=\int_{0}^{R}\int_{0}^{2\pi}\Big(\frac{1}{2\pi}\delta_{y}d\theta\Big)\frac{2r}{R^{2}}~dr=\frac{1}{\pi R^{2}}\iint r~dr~d\theta,$
so we recover the measure $\mu$ by integrating the family $\mu_{y}$ with respect to $\nu$. Moreover, $d(\pi^{-1}(y),\pi^{-1}(y^{\prime}))=|y-y^{\prime}|$. One way to write the 1-Wasserstein distance is
@@
Then, $\mathcal{E}_{1}(\tilde{f})=1$.
[…]
Consider $X=\{(x_{1},x_{2})\in\mathbb{R}^{2}~:~x_{1}^{2}+\lambda^{2}x_{2}^{2}\leq R^{2}\}$, $R>0$, $\lambda\geq 1$, $\mu\in\mathscr{P}(X)$ with full support and uniformly distributed on $X$, $Y=(0,R]$, and a foliation $\mathcal{F}$ of $X$ such that $F_{y}=\{(x_{1},x_{2})\in\mathbb{R}^{2}~:~x_{1}^{2}+\lambda^{2}x_{2}^{2}=y^{2}\}$, for each $y\in Y$, and $\pi:X\to Y$ such that $\pi((x_{1},x_{2}))=\sqrt{x_{1}^{2}+\lambda^{2}x_{2}^{2}}=y$. In polar coordinates
@@
Define $\nu:=\pi_{*}\mu$. Note that, given a open interval $(a,b)\subset(0,R]$, $\nu((a,b))$ is the area between ellipses of major axis $a$ and $b$, normalized by $A$:
@@
Figure 3. Normalized arc length in function of $\theta$ for different $\lambda$.
We have
@@
In this case, $d(\pi^{-1}(y),\pi^{-1}(y^{\prime}))=\frac{1}{\lambda}|y-y^{\prime}|$ and, to calculate the $1$-Wasserstein distance between $\mu_{y}$ and $\mu_{y^{\prime}}$, we still have the radial transport, although $d(x,T(x))$ depends of $\theta$: for $x=y\cos(\theta)$, $d(x,T(x))=R_{y^{\prime}}(\theta)-R_{y}(\theta)$, where
@@
Let $\tilde{f}_{\lambda(t)}$ be the disintegration map of $\mu_{t}$ with respect to $\nu$. Based on the developments above, we have that
$\mathcal{E}_{1}(\tilde{f}_{\lambda(t)})=\frac{2\pi y}{\mathcal{L}_{y}},$
that is, the energy is the ratio between the length of the circle with radius $y$ and the length of the ellipse with major axis equal to $y$. Thus, we have that as $\mu$ evolves over time, the structure for the disintegration of $\mu_{t}$ with respect to $\pi_{*}\mu_{t}$ also changes, and this variation is captured by the energy functional. Indeed, Figure 4 depicts the dependence of $\mathcal{E}_{1}(\tilde{f}_{\lambda(t)})$ on $\lambda(t)$. As shown, $\mathcal{E}_{1}$ grows smoothly with increasing foliation perturbation.
Figure 4. Energy as a function of the eccentricity of the leaves.
For comparison purposes, Table 1 displays the values of $\mathcal{L}_{y}$ for $\lambda=1$ (circle), $\lambda=1,001$, $\lambda=1,01$, $\lambda=1,1$, $\lambda=1,5$ and $\lambda=2$. Table 2 summarizes the computed values of $\mathcal{E}_{1}(\hat{f}_{\lambda})$ corresponding to the cases in Table 1.
@@ 11 $1,000499687$ $1,004968906$ $1,047025412$ $1,188089106$ $1,297046785$
Example 4.6 highlights the versatility of our approach: beyond characterize foliations on $X$ via disintegration maps, it also provides a framework for analyzing structural perturbations in foliations. The concepts developed here have potential applications across Dynamical Systems, Machine Learning, and Stochastic Analysis.
@@ Acknowledgements
Our manuscript has no associated data.
The authors declare no conflict of interest.